# MissionGNN: Hierarchical Multimodal GNN-based Weakly Supervised Video Anomaly Recognition with Mission-Specific Knowledge Graph Generation

### v1

@@ I Introduction
  • Inspired by human intuition, we introduce completely frame-level training and inference that does not require video segmentation which can affect the model’s performance by merging multiple frames into fixed-size segments. Furthermore, this makes our model more practical in real applications by not using future segments or past segments.
@@ II-A Weakly Supervised Video Anomaly Detection and Recognition
The study of anomaly detection in videos has evolved significantly, beginning with methods that heavily relied on fully supervised techniques [2, 42] requiring detailed frame-by-frame labels. To overcome the challenges and high costs associated with detailed annotations, researchers have shifted towards using less resource-intensive approaches. These include weakly supervised methods that train models [17, 36, 38, 48] with broader video-level labels, one-class classification techniques [21, 23, 27] focusing only on normal data, and unsupervised strategies [25, 52] that operate without any labels. Among these, weakly supervised methods have shown notable promise, offering a balance between performance and annotation costs. Our approach aligns with this trend, adopting weakly supervised learning to advance research in this area.
The concept of multiple-instance learning [36] which is dividing videos into segments or bags and treating all frames within a segment as having the same label has become a foundational technique in weakly supervised video anomaly detection. This method, along with enhancements such as internal loss constraints within bags [55] and refinements in segment-level pseudo labeling [57], has facilitated the development of temporal models [38, 17, 48, 53, 44, 57, 39] capable of understanding both short and long-term segment relationships. Unlike these methods, our research seeks new directions beyond the MIL framework to address its limitations, especially in handling variable anomaly durations and the need for real-time analysis which previous works have impracticality in terms of real-time analysis scenarios caused by relying on long-term relationships.
While VAD has received considerable attention, VAR remains underexplored with only a few works focused on it [53], primarily due to the scarcity of samples for each type of anomaly [36]. This scarcity makes VAR a more daunting task than related fields like action recognition [20, 26, 32]. Notably, from the previous work [53], several previous models [38, 47, 17] showed performance drops on some benchmark datasets when it comes to VAR using the mean area under the curve (mAUC) score which is the mean score of the area under the curve (AUC) scores for each anomaly type. However, as Large Language and Vision (LLV) models [29, 31] are introduced to solve action recognition tasks showing success in improving recognition performance [43, 50], an effort to utilize such LLV in VAR tasks is also attempted [53, 54]. These works showed that using LLV that can extract reacher features learned from a large dataset of image-caption pairs can enable enhanced anomaly recognition. Our work introduces an innovative way to adapt large multimodality models to tackle the challenges of VAR in a weakly supervised manner.
@@ II-B Graph Neural Networks and Knowledge Graphs
Graph Neural Networks, including Graph Convolution Networks (GCNs) [56] and Graph Attention Networks (GATs) [41], have been pivotal in processing structural information represented as nodes and edges. The Message Passing Network framework [10] exemplifies a generalized model for GNNs. The network consists of three functions: message function, aggregate function, and update function. Each function can be implemented using simple neural networks with nonlinear functions. Using the message function, it computes messages for all edges and then, each node aggregates messages from neighboring nodes. The aggregation can be sum, mean, or other functions. Finally, it uses the update function to have embeddings of each node from given aggregations. These processes facilitate the generation of node embeddings that capture the essence of the graph’s structure.
GNNs are frequently paired with knowledge graphs to address a wide array of tasks, from question answering [51] and information retrieval [8] to reasoning [9]. Utilizing large-scale, general-domain knowledge graphs like ConceptNet [34], which is one of the widely used semantic knowledge graphs that contain different types of relationships between words, researchers can construct detailed knowledge graphs that enhance model understanding and reasoning capabilities [45, 33, 4]. In our work, we also use ConceptNet in our automated knowledge graph generation framework as one of the structured knowledge sources.
Although a few previous works [57, 39] used GNN as a part of their methods for VAD tasks, their usage of GNN models was limited in capturing temporal relationships between pairs of frames. In contrast, we use a novel GNN architecture over a knowledge graph constructed from our automated knowledge graph generation framework to do reasoning for the VAR task.
@@ II-C Large Multimodal Models and Prompt Optimization
After a huge success in large language models [5, 6, 46], new approaches to have multimodal neural networks [29, 31], which not only can process text but also other forms of modality such as images, emerged. Early models trained to learn joint visual-text embedding spaces with the help of caption-image pairs which are called Large Language and Vision models that can embed given images and text in the same embedding space. LLV models are implemented by having modality-specific encoders and trained with contrastive learning techniques on diverse datasets to align the data representations from different modalities [29, 14]. These LLV models are effective in capturing reacher highly generalized embeddings enabling extension to detection or recognition tasks on images or videos [50, 44, 53] mainly utilizing CLIP [29]. Among those, AnomalyCLIP [53] proposed a model utilizing CLIP to solve the VAR problem demonstrating the strengths of leveraging large multimodal models on VAR tasks by achieving higher mAUC compared to the previous methods. Recently, large multimodal models that can support an expanded range of modalities compared to LLV have been proposed such as ImageBind [11] which can support six different modalities. Our work extends this progress by incorporating ImageBind aiming to explore new possibilities in multimodal anomaly detection although it has similar encoder models with CLIP for text and images.
Prompt optimization techniques have emerged to have better embedding representations to improve models’ performance. Early works conducted human-engineered prompts which manually craft more optimal prompts [15]. Later, to resolve the impracticality of the manual prompting, learnable prompt methods that learn optimal prompt tokens’ embedding using gradients of large multimodal models were introduced [18, 16]. Despite its potential, the learnable prompt approach caused easy overfitting issues [58] with its inefficiency in computing the large model’s gradient hindering its scalability, which makes the approach impractical [61]. Our research proposes an alternative strategy, replacing prompt optimization with knowledge graph reasoning to bypass the limitations of prompt optimization, offering a fresh perspective on enhancing model efficacy.
@@ III-B Mission-Specific Knowledge Graph Generation Framework
Figure 1 shows the overall framework for mission-specific knowledge graph generation. Given a mission $m_{i}$ from a user, the framework generates a KG $G_{m_{i}}=\left(V_{m_{i}},E_{m_{i}}\right)$, which is directed acyclic graph (DAG). In order to generate $G_{m_{i}}$, it first, combines the given mission $m_{i}$ with a pre-formatted prompt to have a query message. Then, $n_{cpt}$ number of key concept words $w^{cpt}_{m_{i},j}$ are extracted by querying the generated query message to a large language model. Each word $w^{cpt}_{m_{i},j}$ is considered as a key concept node $v^{cpt}_{m_{i},j}$ and those key concept nodes build part of the KG $\{v^{cpt}_{m_{i},j}\}=V^{cpt}_{m_{i}}\subset V_{m_{i}}$. Next, to expand KG, iterative sub-graph extraction is conducted using ConceptNet. First sub-graph nodes $v^{sub_{1}}_{m_{i},j}\in V^{sub_{1}}_{m_{i}}$ are retrieved from ConceptNet which are directly neighboring to $v^{cpt}_{m_{i},j}\in V^{cpt}_{m_{i}}$. A set of edges containing neighboring relationships $E^{sub_{1}}_{m_{i}}=\{\left(v,u\right)\in V^{cpt}_{m_{i}}\times V^{sub_{1}}_{m_{i}}|\left(v,u\right)\in\textbf{ConceptNet}\}$ is also extracted during the process. In our implementation, we only considered the “/r/RelatedTo” relationship type. Next sub-graph nodes $v^{sub_{2}}_{m_{i},j}\in V^{sub_{2}}_{m_{i}}$ are extracted with the similar way but directly neighboring to $V^{sub_{1}}_{m_{i}}$ with edges $E^{sub_{2}}_{m_{i}}=\{\left(v,u\right)\in V^{sub_{1}}_{m_{i}}\times V^{sub_{2}}_{m_{i}}|\left(v,u\right)\in\textbf{ConceptNet}\}$. This process, which is extracting $k+1^{th}$ set of sub-graph nodes $V^{sub_{k+1}}_{m_{i}}$ and edges $E^{sub_{k+1}}_{m_{i}}$ from $V^{sub_{k}}_{m_{i}}$, is iteratively conducted until $V^{sub_{d_{sub}}}_{m_{i}}$ and $E^{sub_{d_{sub}}}_{m_{i}}$ are retrieved where $d_{sub}$ indicates maximum depth of the sub-graph. The extracted sub-graph nodes $V^{sub}_{m_{i}}=\bigcup_{1\leq k\leq d_{sub}}{V^{sub_{k}}_{m_{i}}}$ and edges $E^{sub}_{m_{i}}=\bigcup_{1\leq k\leq d_{sub}}{E^{sub_{k}}_{m_{i}}}$ also build part of the KG by $V^{sub}_{m_{i}}\subset V_{m_{i}}$ and $E^{sub}_{m_{i}}\subset E_{m_{i}}$. Finally, $G_{m_{i}}$ is completed by combining sensor node $v^{snr}_{m_{i}}$ and encoding node $v^{ecd}_{m_{i}}$. $v^{snr}_{m_{i}}$ is attached to the graph by connecting it with key concept nodes $v^{cpt}_{m_{i},j}$ towards $v^{cpt}_{m_{i},j}$ direction while $v^{ecd}_{m_{i}}$ is attached by connecting it with leaf nodes, which are nodes that do not have outgoing edges, of sub-graph nodes towards $v^{ecd}_{m_{i}}$ direction. Completed $G_{m_{i}}$ is described in Equation 1.
@@ III-C Hierarchical Multimodal GNN
where $\mathbf{W}^{\phi}_{m_{i},l}$ is trainable parametric matrix, $\mathbf{b}^{\phi}_{m_{i},l}$ is trainable bias, and $X_{l-1}\in\mathbb{R}^{|V_{m_{i}}|\times D_{m_{i},l-1}}$ is embeddings of previous layer for nodes. Hierarchical message passing layer $\mathcal{M}^{h}_{m_{i},l}\colon\mathbb{R}^{|V_{m_{i}}|\times D_{m_{i},l}}\to\mathbb{R}^{|E^{(l)}_{m_{i}}|\times D_{m_{i},l}}$ compute messages each corresponding to each node where $E^{(l)}_{m_{i}}\subset E_{m_{i}}$ indicates a set of edges that connect depth $l-1$ nodes $V^{(l-1)}_{m_{i}}\subset V_{m_{i}}$ and depth $l$ nodes $V^{(l)}_{m_{i}}\subset V_{m_{i}}$. Note that depth 00 node set $V^{(0)}_{m_{i}}$ contains only one node which is sensor node $v^{snr}_{m_{i}}$ and $l$ starts from $1$. Each message indicates measured relativity by the dot product between two knowledge embeddings as described below:
@@
$\begin{split}&\mathcal{G}_{m_{i},l}(X_{l-1})=\\ &\text{ELU}\left(\text{BatchNorm}\left(\mathcal{A}^{h}_{m_{i},l}\left(\phi_{m_{i},l}\left(X_{l-1}\right),{M}^{h}_{m_{i},l}\left(\phi_{m_{i},l}\left(X_{l-1}\right)\right)\right)\right)\right)\end{split}$ (5)
through this, knowledge embeddings of the current layer are computed $X_{l}=\mathcal{G}_{m_{i},l}(X_{l-1})\in\mathbb{R}^{|V_{m_{i}}|\times D_{m_{i},l}}$ from previous layer embeddings $X_{l-1}\in\mathbb{R}^{|V_{m_{i}}|\times D_{m_{i},l-1}}$. It is notable that unlike $\mathcal{M}^{h}_{m_{i},l}$ and $\mathcal{A}^{h}_{m_{i},l}$ where it is conducted only for $V^{(l)}_{m_{i}}$ nodes and $E^{(l)}_{m_{i}}$ edges, other components are applied for all nodes $V_{m_{i}}$ and edges $E_{m_{i}}$ so that embeddings of every node can be placed in the same embedding space for the future reasoning propagation. The final embedding of the KG reasoning $\vec{r}_{T_{i}}$ is extracted from $v^{ecd}_{m_{i}}$ of $X_{d_{sub}+2}$ which is the final embedding layer of KG $G_{T_{i}}$. Finally, reasoning embedding $\vec{f}_{t}$ of the frame image encoding $\mathcal{E}_{I}(F_{t})$ is formed by concatenating all $\vec{r}_{T_{i}}$ from each KG as $\vec{f}_{t}=\vec{r}_{T_{1}}\frown\vec{r}_{T_{2}}\frown\cdots\frown\vec{r}_{T_{n}}\in\mathbb{R}^{D}$ where $D=\sum_{i}{D_{m_{i},d_{sub}+2}}$.
@@ III-E Weakly-supervised Frame-level Training
III-E1 Decaying Threshold-based Anomaly Localization.
A better performing model requires estimating normal frames and anomaly frames among frames in the same video, where a single type of anomaly is labeled in weakly-supervised learning. However, it is difficult to implement a widely used rank-based approach without using MIL-based training. To resolve this issue and have a model that is capable of fully frame-level training and inference, we introduce a novel anomaly localization method, which is a method that separates normal frames from anomaly-labeled video, using our proposed concept of decaying threshold. Decaying threshold $\theta_{d}$ is denoted as follows:
@@
III-E2 Cross-entropy Loss.
After the decaying threshold-based anomaly localization, we have normal frames $F^{N}_{t}$, normal estimated frames $F^{A+}_{t}$, and anomaly estimated frames $F^{A-}_{t}$. We use the cross-entropy loss to maximize probabilities that match each case. Thus, we have three cross-entropy terms for each set of cases as follows:
@@
III-E3 Sparsity and Smoothing Losses.
We adapted the same two regularization loss terms that are used in [37]. First is done through sparsity loss term exploiting that the anomaly cases are rare in entire time frames. Sparsity loss tries to minimize the number of frames that are predicted as anomalies by following:
@@ IV-A Experimental Settings
We perform evaluations that include both VAD and VAR. Consistent with previous research, VAD performance is measured using the area under the curve (AUC) of the receiver operating characteristics (ROC) at the frame level. This metric is chosen because it is independent of thresholding in the detection task. A higher AUC at the frame level indicates a better ability to distinguish normal from abnormal events. For the VAR evaluation, we extend the AUC metric to the multi-classification scenario. Specifically, we compute the AUC for each anomalous class, treating anomalous frames of the class as positive instances and all others as negative instances. We then compute the mean AUC (mAUC) across all anomalous classes. Likewise, for the XD-Violence dataset, we follow the established evaluation protocol [49] and VAD results are presented using the average precision (AP) derived from the precision-recall curve (PRC). For VAR outcomes, we report the mean average precision (mAP), which involves averaging the binary AP values across all abnormal classes.

### v2

@@ I Introduction
  • Inspired by human intuition, we introduce completely frame-level training and inference that does not require video segmentation which can affect the model’s performance by merging multiple frames into fixed-size segments. Furthermore, this makes our model more practical in real applications by not using future segments or past segments suitable for real-time applications.
@@ II-A Weakly Supervised Video Anomaly Detection and Recognition
Video anomaly detection has traditionally relied on fully supervised methods requiring frame-level annotations, which are costly and time-consuming [2, 33]. To address these challenges, weakly supervised approaches using video-level labels [13, 27, 29, 37], one-class classification [16, 18], and unsupervised methods [20, 41] have gained traction. Weakly supervised techniques, in particular, offer a balance between performance and annotation costs, making them popular in recent research.
MIL [27], which divides videos into segments and labels them uniformly, is a key method in weakly supervised video anomaly detection. Enhancements like internal loss constraints [44] and pseudo labeling [46] have enabled models to capture both short- and long-term relationships [29, 13]. However, MIL struggles with variable anomaly durations and is impractical for real-time scenarios due to its reliance on long-term dependencies, which our work seeks to overcome its limitations.
While VAD has been widely studied, VAR is less explored due to the scarcity of labeled anomalies [27]. Recent efforts to apply Large Language and Vision (LLV) models [22, 24] in VAR have shown promising results [42, 43]. These models, trained on large image-caption datasets, can improve anomaly recognition. Our work adapts LLV models for weakly supervised VAR, addressing key challenges in this domain.
@@ II-B Graph Neural Networks and Knowledge Graphs
Graph Neural Networks, such as Graph Convolution Networks (GCNs) [45] and Graph Attention Networks (GATs) [32], are widely used to process structural data represented as nodes and edges. The Message Passing Network framework [7] provides a generalized GNN model, where messages are computed for all edges, aggregated by nodes, and updated to generate node embeddings. These embeddings capture the graph’s structural essence and are critical for tasks requiring an understanding of relationships between entities.
GNNs are often combined with knowledge graphs to enhance tasks like question answering [40], information retrieval [5], and reasoning [6]. A key resource in this space is ConceptNet [25], a large-scale semantic knowledge graph that encodes relationships between words, such as “IsA” or “RelatedTo,” across multiple languages. ConceptNet is widely used in various domains to provide structured context and reasoning capabilities. In our work, ConceptNet serves as one of the core components of our automated knowledge graph generation framework, supplying structured knowledge that aids in anomaly recognition.
While previous approaches [46, 30] have used GNNs for VAD, they mainly focused on capturing temporal relationships between video frames. In contrast, our method employs a novel GNN architecture over a knowledge graph constructed through automated generation to perform semantic reasoning for VAR tasks, extending GNN capabilities beyond temporal analysis.
@@ II-C Large Multimodal Models and Prompt Optimization
Large Language and Vision models [22, 24], which integrate text and image data in a joint embedding space, have proven successful in various tasks, including anomaly detection [39, 35]. These models use modality-specific encoders and contrastive learning to align representations across different modalities. AnomalyCLIP [42], which leverages CLIP [22], demonstrated improved VAR performance by harnessing the strengths of LLV models. Expanding on this, recent models like ImageBind [8] support multiple modalities. Our work builds on this progress by incorporating ImageBind to explore multimodal anomaly detection.
Prompt optimization techniques were introduced to enhance model performance by refining embedding representations. Initial methods relied on manually crafted prompts [11], followed by learnable prompt techniques [14, 12], which automatically optimize prompts using gradients. However, these approaches often face overfitting and scalability issues [47]. To overcome these limitations, our research replaces prompt optimization with knowledge graph reasoning, which acts as knowledge extraction from a large language model, providing a more efficient and scalable approach for improving model efficacy.
@@ III-B Mission-Specific Knowledge Graph Generation Framework
Figure 1 shows the overall framework for mission-specific knowledge graph generation. Given a mission $m_{i}$ from a user, the framework generates a KG $G_{m_{i}}=\left(V_{m_{i}},E_{m_{i}}\right)$, which is directed acyclic graph (DAG). In order to generate $G_{m_{i}}$, it first, combines the given mission $m_{i}$ with a pre-formatted prompt to have a query message. Then, $n_{cpt}$ number of key concept words $w^{cpt}_{m_{i},j}$ are extracted by querying the generated query message to a large language model. Each word $w^{cpt}_{m_{i},j}$ is considered as a key concept node $v^{cpt}_{m_{i},j}$ and those key concept nodes build part of the KG $\{v^{cpt}_{m_{i},j}\}=V^{cpt}_{m_{i}}\subset V_{m_{i}}$. Next, to expand KG, iterative sub-graph extraction is conducted using the large language model. First sub-graph nodes $v^{sub_{1}}_{m_{i},j}\in V^{sub_{1}}_{m_{i}}$ are extracted by querying the key concept nodes to the large language model. A set of edges containing neighboring relationships $E^{sub_{1}}_{m_{i}}=\{\left(v,u\right)\in V^{cpt}_{m_{i}}\times V^{sub_{1}}_{m_{i}}|\left(v,u\right)\in\textbf{ConceptNet}\}$ is also extracted during the process. In our implementation, we only considered the “/r/RelatedTo” relationship type. Next sub-graph nodes $v^{sub_{2}}_{m_{i},j}\in V^{sub_{2}}_{m_{i}}$ are extracted with the similar way but directly neighboring to $V^{sub_{1}}_{m_{i}}$ with edges $E^{sub_{2}}_{m_{i}}=\{\left(v,u\right)\in V^{sub_{1}}_{m_{i}}\times V^{sub_{2}}_{m_{i}}|\left(v,u\right)\in\textbf{ConceptNet}\}$. This process, which is extracting $k+1^{th}$ set of sub-graph nodes $V^{sub_{k+1}}_{m_{i}}$ and edges $E^{sub_{k+1}}_{m_{i}}$ from $V^{sub_{k}}_{m_{i}}$, is iteratively conducted until $V^{sub_{d_{sub}}}_{m_{i}}$ and $E^{sub_{d_{sub}}}_{m_{i}}$ are retrieved where $d_{sub}$ indicates maximum depth of the sub-graph. The extracted sub-graph nodes $V^{sub}_{m_{i}}=\bigcup_{1\leq k\leq d_{sub}}{V^{sub_{k}}_{m_{i}}}$ and edges $E^{sub}_{m_{i}}=\bigcup_{1\leq k\leq d_{sub}}{E^{sub_{k}}_{m_{i}}}$ also build part of the KG by $V^{sub}_{m_{i}}\subset V_{m_{i}}$ and $E^{sub}_{m_{i}}\subset E_{m_{i}}$. Finally, $G_{m_{i}}$ is completed by combining sensor node $v^{snr}_{m_{i}}$ and encoding node $v^{ecd}_{m_{i}}$. $v^{snr}_{m_{i}}$ is attached to the graph by connecting it with key concept nodes $v^{cpt}_{m_{i},j}$ towards $v^{cpt}_{m_{i},j}$ direction while $v^{ecd}_{m_{i}}$ is attached by connecting it with leaf nodes, which are nodes that do not have outgoing edges, of sub-graph nodes towards $v^{ecd}_{m_{i}}$ direction. Completed $G_{m_{i}}$ is described in Equation 1.
@@ III-C Hierarchical Multimodal GNN
where $\mathbf{W}^{\phi}_{m_{i},l}$ is trainable parametric matrix, $\mathbf{b}^{\phi}_{m_{i},l}$ is trainable bias, and $X_{l-1}\in\mathbb{R}^{|V_{m_{i}}|\times D_{m_{i},l-1}}$ is embeddings of previous layer for nodes. Hierarchical message passing layer $\mathcal{M}^{h}_{m_{i},l}\colon\mathbb{R}^{|V_{m_{i}}|\times D_{m_{i},l}}\to\mathbb{R}^{|E^{(l)}_{m_{i}}|\times D_{m_{i},l}}$ compute messages each corresponding to each node where $E^{(l)}_{m_{i}}\subset E_{m_{i}}$ indicates a set of edges that connect depth $l-1$ nodes $V^{(l-1)}_{m_{i}}\subset V_{m_{i}}$ and depth $l$ nodes $V^{(l)}_{m_{i}}\subset V_{m_{i}}$. Note that depth 00 node set $V^{(0)}_{m_{i}}$ contains only one node which is sensor node $v^{snr}_{m_{i}}$ and $l$ starts from $1$. Each message indicates measured relativity by the element-wise product between two knowledge embeddings as described below:
@@
$\begin{split}&\mathcal{G}_{m_{i},l}(X_{l-1})=\text{ELU}(\\ &\text{BatchNorm}\left(\mathcal{A}^{h}_{m_{i},l}\left(\phi_{m_{i},l}\left(X_{l-1}\right),{M}^{h}_{m_{i},l}\left(\phi_{m_{i},l}\left(X_{l-1}\right)\right)\right)\right))\end{split}$ (5)
[…]
@@ III-E Weakly-supervised Frame-level Training
Decaying Threshold-based Anomaly Localization. A better performing model requires estimating normal frames and anomaly frames among frames in the same video, where a single type of anomaly is labeled in weakly-supervised learning. However, it is difficult to implement a widely used rank-based approach without using MIL-based training. To resolve this issue and have a model that is capable of fully frame-level training and inference, we introduce a novel anomaly localization method, which is a method that separates normal frames from anomaly-labeled video, using our proposed concept of decaying threshold. Decaying threshold $\theta_{d}$ is denoted as follows:
@@
Cross-entropy Loss. After the decaying threshold-based anomaly localization, we have normal frames $F^{N}_{t}$, normal estimated frames $F^{A+}_{t}$, and anomaly estimated frames $F^{A-}_{t}$. We use the cross-entropy loss to maximize probabilities that match each case. Thus, we have three cross-entropy terms for each set of cases as follows:
@@
Sparsity and Smoothing Losses. We adapted the same two regularization loss terms that are used in [28]. First is done through sparsity loss term exploiting that the anomaly cases are rare in entire time frames. Sparsity loss tries to minimize the number of frames that are predicted as anomalies by following:
@@ IV-A Experimental Settings
We perform evaluations that include both VAD and VAR. Consistent with previous research, VAD performance is measured using the area under the curve (AUC) of the receiver operating characteristics (ROC) at the frame level. A higher AUC at the frame level indicates a better ability to distinguish normal from abnormal events. For the VAR evaluation, we extend the AUC metric to the multi-classification scenario. Specifically, we compute the AUC for each anomalous class, treating anomalous frames of the class as positive instances and all others as negative instances. We then compute the mean AUC (mAUC) across all anomalous classes. Likewise, for the XD-Violence dataset, we follow the established evaluation protocol [38] and VAD results are presented using the average precision (AP) derived from the precision-recall curve (PRC). For VAR outcomes, we report the mean average precision (mAP), which involves averaging the binary AP values across all abnormal classes.